ADNAN MUHAMMED

Grade 12,

a stone is dropped from the top of a building .how long does it Take to fall 20 m? how fast does it move at the end of the fall .how time taken before it reaches the ground

a stone is dropped from the top of a building .how long does it Take to fall 20 m? how fast does it move at the end of the fall .how time taken before it reaches the ground

Grade:12th pass

2 Answers

Arun
6 years ago
Dear student
 
Using
S = ut + ½ * g* t²
20 = 0 + ½ * 10 * t²
t² = 40/10 = 4
t = 2 sec.
 
Regards
Arun (askIITians forum expert)
 
 
Khimraj
6 years ago
You can solve by use of equation of motion.
S = ut + ½gt2
20 = 0 + 5t2
t = 2 sec
v2 = u2 + 2gs
v2 = 0+ 2*10*20
v = 20m/s.
Hope it clears. If you like answer then please approve it.